# Detection of Gravitational Anomaly at Low Acceleration from a Highest-quality Sample of 36 Wide Binaries with Accurate 3D Velocities

### v1

@@ E.5 Speckle Holography
Gaia DR3 1003223584897948160 2024.8650 0.28 0.20
Gaia DR3 1003223614961194752 2024.8650 0.38 0.27
Gaia DR3 1020666649676726784 2025.2123 0.26 0.20
Gaia DR3 1020667714828594304 2025.2123 0.37 0.26
Gaia DR3 1023887978228642176 2025.2176 0.31 0.22
Gaia DR3 1023888012588380032 2024.8680 0.25 0.18
Gaia DR3 1028432805245868544 2024.8680 0.24 0.17
Gaia DR3 1028432805250731264 2024.8680 0.26 0.19
Gaia DR3 104998928046427264 2024.6869 0.32 0.22
Gaia DR3 105004842216905344 2024.6869 0.32 0.22
Gaia DR3 1159502154714284544 2025.4422 1.00 0.61
Gaia DR3 1159502498311670144 2025.4422 0.19 0.14
Gaia DR3 1172915990414659328 2025.4393 0.18 0.12
Gaia DR3 1172920487244742912 2025.4393 0.19 0.13
Gaia DR3 1174143182830505856 2025.4448 0.17 0.12
Gaia DR3 1174143182830505984 2025.4448 0.11 0.10
Gaia DR3 1178121422057906688 2025.4393 0.19 0.13
Gaia DR3 1178121422057907200 2025.4393 0.27 0.19
Gaia DR3 1182257922240686720 2025.4339 1.33 1.01
Gaia DR3 1182257922240687232 2025.4339 1.27 0.90
Gaia DR3 1214182826468995584 2025.4367 0.45 0.30
Gaia DR3 1214182826468995840 2025.4367 0.55 0.34
Gaia DR3 121690893241571968 2024.8618 0.92 0.49
Gaia DR3 121691266902954240 2024.8618 0.64 0.37
Gaia DR3 1224242120913963776 2025.4340 0.40 0.27
Gaia DR3 1224242120913963904 2025.4340 0.39 0.25
Gaia DR3 1227416514060783232 2025.4448 0.45 0.31
Gaia DR3 1227417304334766208 2025.4366 0.27 0.21
Gaia DR3 1247610934890765184 2025.4420 0.53 0.36
Gaia DR3 1247610934890765312 2025.4420 0.39 0.26
Gaia DR3 1258410612976538368 2025.4393 0.10 0.09
Gaia DR3 1258410750415492864 2025.4393 0.13 0.11
Gaia DR3 1271824002003198464 2025.4367 0.98 0.59
Gaia DR3 1271824002003198976 2025.4367 0.98 0.58
Gaia DR3 1274562370068531712 2025.4367 0.19 0.13
Gaia DR3 1274568245587206144 2025.4367 0.13 0.11
Gaia DR3 1282815063829295360 2025.4393 0.10 0.09
Gaia DR3 1282817022334383104 2025.4394 0.11 0.10
Gaia DR3 1282817022334383232 2025.2182 0.18 0.13
Gaia DR3 1297419700686304896 2025.4341 0.84 0.45
Gaia DR3 1297419808063095936 2025.4342 0.71 0.40
Gaia DR3 1299746168507031424 2025.4450 0.17 0.12
Gaia DR3 1299746168508720384 2025.4450 0.17 0.12
Gaia DR3 1302449489644570496 2025.4395 0.00 0.00
Gaia DR3 1302543326089734144 2025.4395 1.07 0.65
Gaia DR3 1305035575352515840 2025.4396 0.50 0.32
Gaia DR3 1305036228187622912 2025.4396 0.00 0.00
Gaia DR3 1306220264772422016 2025.4450 0.29 0.21
Gaia DR3 1306220642729543936 2025.4450 0.28 0.20
Gaia DR3 1308215328620272256 2025.4450 0.28 0.20
Gaia DR3 1308215328620272384 2025.4450 0.26 0.19
Gaia DR3 1320009686772023296 2025.4367 0.34 0.23
Gaia DR3 1320009721131760640 2025.4367 0.27 0.19
Gaia DR3 1320379809873587968 2025.4396 0.38 0.27
Gaia DR3 1321677920789034752 2025.4341 0.45 0.30
Gaia DR3 1321677920789214080 2025.4341 0.69 0.42
Gaia DR3 1327463825851820800 2025.4450 0.28 0.19
Gaia DR3 1327463825851820928 2025.4450 0.26 0.18
Gaia DR3 1331462302965590144 2025.4396 0.00 0.00
Gaia DR3 1331462302965590400 2025.4396 0.19 0.13
Gaia DR3 1332438183958298496 2025.4423 0.40 0.28
Gaia DR3 1332438188254551808 2025.4423 0.49 0.33
Gaia DR3 1332473922382506624 2025.4341 1.32 1.01
Gaia DR3 1332473922382951424 2025.4341 1.02 0.57
Gaia DR3 1346694112423491712 2025.4370 0.91 0.53
Gaia DR3 1346694146783637504 2025.4370 0.30 0.22
Gaia DR3 1348285896022947584 2025.4370 0.13 0.11
Gaia DR3 1348286651937191040 2025.4370 0.23 0.16
Gaia DR3 1352381027080312832 2025.4423 0.19 0.14
Gaia DR3 1352427996842670464 2025.4423 0.58 0.38
Gaia DR3 1356763852227145344 2025.4424 1.03 0.73
Gaia DR3 1356763886586884864 2025.4424 0.92 0.59
Gaia DR3 1357040345041982464 2025.4424 0.90 0.58
Gaia DR3 1357040379401720448 2025.4424 0.44 0.31
Gaia DR3 1371903199628510464 2025.4394 0.15 0.12
Gaia DR3 1371903199628531200 2025.4394 0.21 0.13
Gaia DR3 1379737421839743104 2025.4341 0.30 0.21
Gaia DR3 1379737426135359360 2025.4341 0.29 0.20
Gaia DR3 1381408782592655616 2025.4424 0.45 0.31
Gaia DR3 1381455722290898176 2025.4424 1.13 0.80
Gaia DR3 1391176092275606528 2025.4394 0.00 0.00
Gaia DR3 1391176092275606656 2025.4394 0.29 0.20
Gaia DR3 1396388773823564416 2025.4340 0.34 0.24
Gaia DR3 1396388773823907584 2025.4340 0.31 0.22
Gaia DR3 1410982832177435008 2025.4341 1.13 0.72
Gaia DR3 1410982862240353664 2025.4341 0.40 0.27
Gaia DR3 1419999819661235840 2025.4344 0.19 0.13
Gaia DR3 1420000579871414528 2025.4344 0.17 0.12
Gaia DR3 1434064639260540288 2025.4397 0.37 0.27
Gaia DR3 1434064669325509760 2025.4397 1.15 0.78
Gaia DR3 1444305829863259648 2025.4420 0.54 0.36
Gaia DR3 1444305937237949184 2025.4420 0.92 0.60
Gaia DR3 1467776078085413888 2025.4392 0.31 0.22
Gaia DR3 1467776078085414400 2025.4393 0.32 0.23
Gaia DR3 1478999686625786368 2025.4366 0.28 0.21
Gaia DR3 1478999755345263744 2025.4366 0.39 0.28
Gaia DR3 1479430076707007488 2025.4394 0.25 0.17
Gaia DR3 1482432155767129728 2025.4394 0.13 0.11
Gaia DR3 1508470171660833536 2025.4366 0.53 0.33
Gaia DR3 1508470206020209664 2025.4366 0.53 0.34
Gaia DR3 1509241238549271808 2025.4365 0.34 0.25
Gaia DR3 1509241238549292800 2025.4365 0.62 0.39
Gaia DR3 1509783778818258176 2025.4365 0.23 0.18
Gaia DR3 1509783778818258304 2025.4365 0.33 0.25
Gaia DR3 1527328376624427776 2025.4393 0.33 0.24
Gaia DR3 1527328376624971520 2025.4393 0.35 0.26
Gaia DR3 153741691551129216 2024.1277 0.40 0.28
Gaia DR3 153741760270606464 2024.1277 0.40 0.28
Gaia DR3 1543083553620064768 2024.1286 0.00 0.00
Gaia DR3 1543083691059006592 2024.1286 0.38 0.28
Gaia DR3 1549520949383005568 2025.4365 0.28 0.21
Gaia DR3 1549521125478552064 2025.4365 0.18 0.13
Gaia DR3 1585800782946765312 2025.4448 0.46 0.32
Gaia DR3 1585800851666242688 2025.4448 0.25 0.19
Gaia DR3 1586977737129182848 2025.4448 0.20 0.14
Gaia DR3 1586977844504488576 2025.4448 0.13 0.11
Gaia DR3 1590641417247939328 2025.4394 0.00 0.00
Gaia DR3 1590641417247939584 2025.4394 0.82 0.52
Gaia DR3 1592707502675396992 2025.4339 0.38 0.26
Gaia DR3 1592707502675731456 2025.4339 0.37 0.26
Gaia DR3 1595968929041676544 2025.4340 0.48 0.32
Gaia DR3 1595968929041676672 2025.4340 0.44 0.30
Gaia DR3 1605833880509338752 2025.4448 0.29 0.21
Gaia DR3 1605833884804824704 2025.4448 0.40 0.28
Gaia DR3 1610186095424889088 2025.4448 0.21 0.16
Gaia DR3 1610191970940150272 2025.4448 0.27 0.20
Gaia DR3 1610236260642857344 2025.4448 0.32 0.24
Gaia DR3 1610236260642871552 2025.4448 0.32 0.24
Gaia DR3 162781051522771072 2024.8621 0.38 0.26
Gaia DR3 162781051522771200 2024.8621 0.41 0.29
Gaia DR3 167549908330274176 2024.8618 0.64 0.36
Gaia DR3 167549942690010368 2024.8618 0.94 0.54
Gaia DR3 1730273101846974080 2024.6838 0.54 0.34
Gaia DR3 1730273106143055488 2024.6838 0.44 0.31
Gaia DR3 1738269854212128512 2024.6810 0.35 0.24
Gaia DR3 1738269854213155712 2024.6810 0.37 0.26
Gaia DR3 1749987585851186816 2024.6783 0.57 0.34
Gaia DR3 1749987585854235776 2024.6783 1.84 1.33
Gaia DR3 1760471948915107200 2025.4455 0.18 0.13
Gaia DR3 1760477618271932672 2025.4455 0.26 0.18
Gaia DR3 1762461309047562240 2024.6782 0.45 0.30
Gaia DR3 1762461309047562368 2025.4400 0.35 0.24
Gaia DR3 1762461893163118464 2025.4455 0.33 0.23
Gaia DR3 1762461893163118592 2024.6782 0.67 0.37
Gaia DR3 1778929480673414784 2024.6838 0.23 0.16
Gaia DR3 1778930240883745536 2024.6839 0.00 0.00
Gaia DR3 1781495706454162944 2024.6866 0.27 0.18
Gaia DR3 1781495710748864256 2024.6866 0.23 0.16
Gaia DR3 1789024749766697088 2024.6810 1.16 0.85
Gaia DR3 1789024857138962048 2024.6810 0.20 0.14
The third component is DR3 181318611409243136, Gaia DR3 181312735893067008 2024.8650 0.20 0.15
@@
The third component is DR3 181318611409243136, Gaia DR3 181318611408327552 2024.8650 0.17 0.12
@@
The third component, DR3 1815165879238575232, Gaia DR3 1815165535636339072 2025.4400 0.16 0.12
@@
Gaia DR3 1861752775327557504 2024.6781 0.37 0.25
Gaia DR3 1861752775327558656 2024.6781 0.45 0.30
Gaia DR3 1868257722327168768 2024.6811 0.72 0.38
Gaia DR3 1868257726634854784 2024.6811 0.67 0.36
Gaia DR3 1868985324150840192 2024.6864 0.22 0.14
Gaia DR3 1868985328454953856 2024.6864 0.24 0.15
Gaia DR3 1871558941576158464 2024.6810 0.38 0.25
Gaia DR3 1871559697490418816 2025.4455 0.46 0.30
Gaia DR3 1894003238061739648 2024.6864 1.01 0.61
Gaia DR3 1894003238061739776 2024.6864 1.19 0.74
Gaia DR3 1895940302674258304 2024.6865 1.26 0.82
Gaia DR3 1895940302674258688 2024.6865 0.28 0.19
Gaia DR3 1899383114098895616 2024.6864 0.28 0.19
Gaia DR3 1899383114098895872 2024.6864 0.33 0.22
Gaia DR3 1902676117063909888 2024.6865 0.43 0.29
Gaia DR3 1902679033343158528 2024.6865 0.43 0.28
Gaia DR3 1904247284819944448 2024.6865 0.35 0.25
Gaia DR3 1904247284820463104 2024.6865 0.25 0.17
Gaia DR3 1909810023381344768 2024.6866 0.40 0.27
Gaia DR3 1909810126460559616 2024.6866 0.43 0.29
Gaia DR3 1910071432270904832 2024.6866 0.37 0.25
Gaia DR3 1910071535350120192 2024.6866 0.42 0.28
Gaia DR3 1921150828692050304 2024.6866 0.51 0.34
Gaia DR3 1921150832986231680 2024.6866 1.19 0.79
Gaia DR3 1923116072583279232 2024.6865 0.24 0.16
Gaia DR3 1923116072583279488 2024.6759 0.60 0.39
Gaia DR3 1931673464205703936 2024.6865 0.55 0.34
Gaia DR3 1931673532926948864 2024.6865 0.96 0.52
Gaia DR3 1932297578791924992 2024.6865 0.22 0.14
Gaia DR3 1932297578791925760 2024.6865 0.22 0.14
Gaia DR3 1938247517245907456 2024.6866 0.17 0.12
Gaia DR3 1938247654684985344 2024.6866 0.15 0.12
Gaia DR3 1942018494236897792 2024.6866 0.62 0.36
Gaia DR3 1942018498533031808 2024.6866 0.56 0.34
Gaia DR3 1942384773344557056 2024.6865 0.20 0.13
Gaia DR3 1942384872124424704 2024.6865 0.32 0.22
Gaia DR3 1960631100090155008 2024.6838 0.11 0.10
Gaia DR3 1960631100090155264 2024.6838 0.27 0.20
Gaia DR3 1967282939282261120 2025.4455 0.22 0.15
Gaia DR3 1967283042361454848 2025.4455 0.13 0.11
Gaia DR3 1971147791380289536 2024.6837 0.23 0.16
Gaia DR3 1971148100617936128 2024.6837 0.00 0.00
Gaia DR3 1985752226366174720 2024.6865 0.49 0.33
Gaia DR3 1985752226366174848 2024.6865 0.25 0.17
Gaia DR3 1994221730075916800 2024.6759 0.88 0.58
Gaia DR3 1994221936234345344 2024.6759 1.35 1.07
Gaia DR3 200266181758708736 2024.8649 1.06 0.65
Gaia DR3 200266250476976768 2024.8649 0.36 0.25
The third component, DR3 1815165879238575232, Gaia DR3 2005665584549011328 2025.4400 0.21 0.15
@@
The third component, DR3 2005665584530878848, Gaia DR3 2005665584549011328 2024.6867 0.33 0.23
@@
The third component, DR3 2005665584530878848, Gaia DR3 2005665962510573824 2024.6867 0.40 0.28

### v2

@@ E.5 Speckle Holography
Gaia DR3 1003223584897948160 2024.865 0.140 0.100
Gaia DR3 1003223614961194752 2024.865 0.190 0.135
Gaia DR3 1020666649676726784 2025.2123 0.130 0.100
Gaia DR3 1020667714828594304 2025.2123 0.185 0.130
Gaia DR3 1023887978228642176 2025.2176 0.155 0.110
Gaia DR3 1023888012588380032 2024.868 0.125 0.090
Gaia DR3 1028432805245868544 2024.868 0.120 0.085
Gaia DR3 1028432805250731264 2024.868 0.130 0.095
Gaia DR3 104998928046427264 2024.6869 0.160 0.110
Gaia DR3 105004842216905344 2024.6869 0.160 0.110
Gaia DR3 1159502154714284544 2025.4422 0.500 0.305
Gaia DR3 1159502498311670144 2025.4422 0.095 0.070
Gaia DR3 1172915990414659328 2025.4393 0.090 0.060
Gaia DR3 1172920487244742912 2025.4393 0.095 0.065
Gaia DR3 1174143182830505856 2025.4448 0.085 0.060
Gaia DR3 1174143182830505984 2025.4448 0.055 0.050
Gaia DR3 1178121422057906688 2025.4393 0.095 0.065
Gaia DR3 1178121422057907200 2025.4393 0.135 0.095
Gaia DR3 1182257922240686720 2025.4339 0.665 0.505
Gaia DR3 1182257922240687232 2025.4339 0.635 0.450
Gaia DR3 1214182826468995584 2025.4367 0.225 0.150
Gaia DR3 1214182826468995840 2025.4367 0.275 0.170
Gaia DR3 121690893241571968 2024.8618 0.460 0.245
Gaia DR3 121691266902954240 2024.8618 0.320 0.185
Gaia DR3 1224242120913963776 2025.434 0.200 0.135
Gaia DR3 1224242120913963904 2025.434 0.195 0.125
Gaia DR3 1227416514060783232 2025.4448 0.225 0.155
Gaia DR3 1227417304334766208 2025.4366 0.135 0.105
Gaia DR3 1247610934890765184 2025.442 0.265 0.180
Gaia DR3 1247610934890765312 2025.442 0.195 0.130
Gaia DR3 1258410612976538368 2025.4393 0.050 0.045
Gaia DR3 1258410750415492864 2025.4393 0.065 0.055
Gaia DR3 1271824002003198464 2025.4367 0.490 0.295
Gaia DR3 1271824002003198976 2025.4367 0.490 0.290
Gaia DR3 1274562370068531712 2025.4367 0.095 0.065
Gaia DR3 1274568245587206144 2025.4367 0.065 0.055
Gaia DR3 1282815063829295360 2025.4393 0.050 0.045
Gaia DR3 1282817022334383104 2025.4394 0.055 0.050
Gaia DR3 1282817022334383232 2025.2182 0.090 0.065
Gaia DR3 1297419700686304896 2025.4341 0.420 0.225
Gaia DR3 1297419808063095936 2025.4342 0.355 0.200
Gaia DR3 1299746168507031424 2025.445 0.085 0.060
Gaia DR3 1299746168508720384 2025.445 0.085 0.060
Too weak Gaia DR3 1302449489644570496 2025.4395 >2.>2. >2.>2.
Gaia DR3 1302543326089734144 2025.4395 0.535 0.325
Gaia DR3 1305035575352515840 2025.4396 0.250 0.160
Too weak Gaia DR3 1305036228187622912 2025.4396 >2.>2. >2.>2.
Gaia DR3 1306220264772422016 2025.445 0.145 0.105
Gaia DR3 1306220642729543936 2025.445 0.140 0.100
Gaia DR3 1308215328620272256 2025.445 0.140 0.100
Gaia DR3 1308215328620272384 2025.445 0.130 0.095
Gaia DR3 1320009686772023296 2025.4367 0.170 0.115
Gaia DR3 1320009721131760640 2025.4367 0.135 0.095
Gaia DR3 1320379809873587968 2025.4396 0.190 0.135
Gaia DR3 1321677920789034752 2025.4341 0.225 0.150
Gaia DR3 1321677920789214080 2025.4341 0.345 0.210
Gaia DR3 1327463825851820800 2025.445 0.140 0.095
Gaia DR3 1327463825851820928 2025.445 0.130 0.090
Too weak Gaia DR3 1331462302965590144 2025.4396 >2.>2. >2.>2.
Gaia DR3 1331462302965590400 2025.4396 0.095 0.065
Gaia DR3 1332438183958298496 2025.4423 0.200 0.140
Gaia DR3 1332438188254551808 2025.4423 0.245 0.165
Gaia DR3 1332473922382506624 2025.4341 0.660 0.505
Gaia DR3 1332473922382951424 2025.4341 0.510 0.285
Gaia DR3 1346694112423491712 2025.437 0.455 0.265
Gaia DR3 1346694146783637504 2025.437 0.150 0.110
Gaia DR3 1348285896022947584 2025.437 0.065 0.055
Gaia DR3 1348286651937191040 2025.437 0.115 0.080
Gaia DR3 1352381027080312832 2025.4423 0.095 0.070
Gaia DR3 1352427996842670464 2025.4423 0.290 0.190
Gaia DR3 1356763852227145344 2025.4424 0.515 0.365
Gaia DR3 1356763886586884864 2025.4424 0.460 0.295
Gaia DR3 1357040345041982464 2025.4424 0.450 0.290
Gaia DR3 1357040379401720448 2025.4424 0.220 0.155
Gaia DR3 1371903199628510464 2025.4394 0.075 0.060
Gaia DR3 1371903199628531200 2025.4394 0.105 0.065
Gaia DR3 1379737421839743104 2025.4341 0.150 0.105
Gaia DR3 1379737426135359360 2025.4341 0.145 0.100
Gaia DR3 1381408782592655616 2025.4424 0.225 0.155
Gaia DR3 1381455722290898176 2025.4424 0.565 0.400
Too weak Gaia DR3 1391176092275606528 2025.4394 >2.>2. >2.>2.
Gaia DR3 1391176092275606656 2025.4394 0.145 0.100
Gaia DR3 1396388773823564416 2025.434 0.170 0.120
Gaia DR3 1396388773823907584 2025.434 0.155 0.110
Gaia DR3 1410982832177435008 2025.4341 0.565 0.360
Gaia DR3 1410982862240353664 2025.4341 0.200 0.135
Gaia DR3 1419999819661235840 2025.4344 0.095 0.065
Gaia DR3 1420000579871414528 2025.4344 0.085 0.060
Gaia DR3 1434064639260540288 2025.4397 0.185 0.135
Gaia DR3 1434064669325509760 2025.4397 0.575 0.390
Gaia DR3 1444305829863259648 2025.442 0.270 0.180
Gaia DR3 1444305937237949184 2025.442 0.460 0.300
Gaia DR3 1467776078085413888 2025.4392 0.155 0.110
Gaia DR3 1467776078085414400 2025.4393 0.160 0.115
Gaia DR3 1478999686625786368 2025.4366 0.140 0.105
Gaia DR3 1478999755345263744 2025.4366 0.195 0.140
Gaia DR3 1479430076707007488 2025.4394 0.125 0.085
Gaia DR3 1482432155767129728 2025.4394 0.065 0.055
Gaia DR3 1508470171660833536 2025.4366 0.265 0.165
Gaia DR3 1508470206020209664 2025.4366 0.265 0.170
Gaia DR3 1509241238549271808 2025.4365 0.170 0.125
Gaia DR3 1509241238549292800 2025.4365 0.310 0.195
Gaia DR3 1509783778818258176 2025.4365 0.115 0.090
Gaia DR3 1509783778818258304 2025.4365 0.165 0.125
Gaia DR3 1527328376624427776 2025.4393 0.165 0.120
Gaia DR3 1527328376624971520 2025.4393 0.175 0.130
Gaia DR3 153741691551129216 2024.1277 0.200 0.140
Gaia DR3 153741760270606464 2024.1277 0.200 0.140
Too weak Gaia DR3 1543083553620064768 2024.1286 >2.>2. >2.>2.
Gaia DR3 1543083691059006592 2024.1286 0.190 0.140
Gaia DR3 1549520949383005568 2025.4365 0.140 0.105
Gaia DR3 1549521125478552064 2025.4365 0.090 0.065
Gaia DR3 1585800782946765312 2025.4448 0.230 0.160
Gaia DR3 1585800851666242688 2025.4448 0.125 0.095
Gaia DR3 1586977737129182848 2025.4448 0.100 0.070
Gaia DR3 1586977844504488576 2025.4448 0.065 0.055
Too weak Gaia DR3 1590641417247939328 2025.4394 >2.>2. >2.>2.
Gaia DR3 1590641417247939584 2025.4394 0.410 0.260
Gaia DR3 1592707502675396992 2025.4339 0.190 0.130
Gaia DR3 1592707502675731456 2025.4339 0.185 0.130
Gaia DR3 1595968929041676544 2025.434 0.240 0.160
Gaia DR3 1595968929041676672 2025.434 0.220 0.150
Gaia DR3 1605833880509338752 2025.4448 0.145 0.105
Gaia DR3 1605833884804824704 2025.4448 0.200 0.140
Gaia DR3 1610186095424889088 2025.4448 0.105 0.080
Gaia DR3 1610191970940150272 2025.4448 0.135 0.100
Gaia DR3 1610236260642857344 2025.4448 0.160 0.120
Gaia DR3 1610236260642871552 2025.4448 0.160 0.120
Gaia DR3 162781051522771072 2024.8621 0.190 0.130
Gaia DR3 162781051522771200 2024.8621 0.205 0.145
Gaia DR3 167549908330274176 2024.8618 0.320 0.180
Gaia DR3 167549942690010368 2024.8618 0.470 0.270
Gaia DR3 1730273101846974080 2024.6838 0.270 0.170
Gaia DR3 1730273106143055488 2024.6838 0.220 0.155
Gaia DR3 1738269854212128512 2024.681 0.175 0.120
Gaia DR3 1738269854213155712 2024.681 0.185 0.130
Gaia DR3 1749987585851186816 2024.6783 0.285 0.170
Gaia DR3 1749987585854235776 2024.6783 0.920 0.665
Gaia DR3 1760471948915107200 2025.4455 0.090 0.065
Gaia DR3 1760477618271932672 2025.4455 0.130 0.090
Gaia DR3 1762461309047562240 2024.6782 0.225 0.150
Gaia DR3 1762461309047562368 2025.44 0.175 0.120
Gaia DR3 1762461893163118464 2025.4455 0.165 0.115
Gaia DR3 1762461893163118592 2024.6782 0.335 0.185
Gaia DR3 1778929480673414784 2024.6838 0.115 0.080
Too weak Gaia DR3 1778930240883745536 2024.6839 >2.>2. >2.>2.
Gaia DR3 1781495706454162944 2024.6866 0.135 0.090
Gaia DR3 1781495710748864256 2024.6866 0.115 0.080
Gaia DR3 1789024749766697088 2024.681 0.580 0.425
Gaia DR3 1789024857138962048 2024.681 0.100 0.070
The third component is DR3 181318611409243136, Gaia DR3 181312735893067008 2024.865 0.100 0.075
@@
The third component is DR3 181318611409243136, Gaia DR3 181318611408327552 2024.865 0.085 0.060
@@
The third component, DR3 1815165879238575232, Gaia DR3 1815165535636339072 2025.44 0.080 0.060
@@
Gaia DR3 1861752775327557504 2024.6781 0.185 0.125
Gaia DR3 1861752775327558656 2024.6781 0.225 0.150
Gaia DR3 1868257722327168768 2024.6811 0.360 0.190
Gaia DR3 1868257726634854784 2024.6811 0.335 0.180
Gaia DR3 1868985324150840192 2024.6864 0.110 0.070
Gaia DR3 1868985328454953856 2024.6864 0.120 0.075
Gaia DR3 1871558941576158464 2024.681 0.190 0.125
Gaia DR3 1871559697490418816 2025.4455 0.230 0.150
Gaia DR3 1894003238061739648 2024.6864 0.505 0.305
Gaia DR3 1894003238061739776 2024.6864 0.595 0.370
Gaia DR3 1895940302674258304 2024.6865 0.630 0.410
Gaia DR3 1895940302674258688 2024.6865 0.140 0.095
Gaia DR3 1899383114098895616 2024.6864 0.140 0.095
Gaia DR3 1899383114098895872 2024.6864 0.165 0.110
Gaia DR3 1902676117063909888 2024.6865 0.215 0.145
Gaia DR3 1902679033343158528 2024.6865 0.215 0.140
Gaia DR3 1904247284819944448 2024.6865 0.175 0.125
Gaia DR3 1904247284820463104 2024.6865 0.125 0.085
Gaia DR3 1909810023381344768 2024.6866 0.200 0.135
Gaia DR3 1909810126460559616 2024.6866 0.215 0.145
Gaia DR3 1910071432270904832 2024.6866 0.185 0.125
Gaia DR3 1910071535350120192 2024.6866 0.210 0.140
Gaia DR3 1921150828692050304 2024.6866 0.255 0.170
Gaia DR3 1921150832986231680 2024.6866 0.595 0.395
Gaia DR3 1923116072583279232 2024.6865 0.120 0.080
Gaia DR3 1923116072583279488 2024.6759 0.300 0.195
Gaia DR3 1931673464205703936 2024.6865 0.275 0.170
Gaia DR3 1931673532926948864 2024.6865 0.480 0.260
Gaia DR3 1932297578791924992 2024.6865 0.110 0.070
Gaia DR3 1932297578791925760 2024.6865 0.110 0.070
Gaia DR3 1938247517245907456 2024.6866 0.085 0.060
Gaia DR3 1938247654684985344 2024.6866 0.075 0.060
Gaia DR3 1942018494236897792 2024.6866 0.310 0.180
Gaia DR3 1942018498533031808 2024.6866 0.280 0.170
Gaia DR3 1942384773344557056 2024.6865 0.100 0.065
Gaia DR3 1942384872124424704 2024.6865 0.160 0.110
Gaia DR3 1960631100090155008 2024.6838 0.055 0.050
Gaia DR3 1960631100090155264 2024.6838 0.135 0.100
Gaia DR3 1967282939282261120 2025.4455 0.110 0.075
Gaia DR3 1967283042361454848 2025.4455 0.065 0.055
Gaia DR3 1971147791380289536 2024.6837 0.115 0.080
Too weak Gaia DR3 1971148100617936128 2024.6837 >2.>2. >2.>2.
Gaia DR3 1985752226366174720 2024.6865 0.245 0.165
Gaia DR3 1985752226366174848 2024.6865 0.125 0.085
Gaia DR3 1994221730075916800 2024.6759 0.440 0.290
Gaia DR3 1994221936234345344 2024.6759 0.675 0.535
Gaia DR3 200266181758708736 2024.8649 0.530 0.325
Gaia DR3 200266250476976768 2024.8649 0.180 0.125
The third component, DR3 1815165879238575232, Gaia DR3 2005665584549011328 2025.44 0.105 0.075
@@
The third component, DR3 2005665584530878848, Gaia DR3 2005665584549011328 2024.6867 0.165 0.115
@@
The third component, DR3 2005665584530878848, Gaia DR3 2005665962510573824 2024.6867 0.200 0.140